Part A An ideal gas expands through an adiabatic process. Which of the following statements is/are true? Check all that apply. Check all that apply.

The work done by the gas is negative, and heat must be added to the system.

The work done by the gas is positive, and no heat exchange occurs. The internal energy of the system has increased.

The internal energy of the system has decreased. SubmitHintsMy AnswersGive UpReview Part Incorrect; Try Again; 5 attempts remaining

You are on the right track, but keep in mind that in an adiabatic process Q=0 and ΔU=−W. If the work done is positive, does the internal energy increase or decrease? Can someone explain to me what is going on. I am confuse about thermodynamics principle in the question
Answer #1
Concepts and reason

The main concepts used to solve this problem are first law of thermodynamics and adiabatic process.

Initially, apply the first law of thermodynamics. Later, use the fact that when a gas expands it does work on its surroundings. Finally, use the adiabatic process definition to find the correct statement(s).

Fundamentals

The first law of thermodynamics is expressed as follows:

dQ=dU+dWdQ = dU + dW

Here, dQdQ is the amount of heat added, dU is the change in internal energy, and dW is the work extracted from the system.

An adiabatic process is a thermodynamic process in which no heat exchange occurs.

dQ=0dQ = 0

(A.1)

The first law of thermodynamics is expressed as follows:

dQ=dU+dWdQ = dU + dW

Here, dQdQ is the amount of heat added, dU is the change in internal energy, and dW is the work extracted from the system.

Substitute 0 for dQ in expression dQ=dU+dWdQ = dU + dW.

0=dU+dWdU=dW\begin{array}{c}\\0 = dU + dW\\\\dU = - dW\\\end{array}

It means the internal energy is negative so it decreases.

(A.2)

An adiabatic process is a thermodynamic process in which no heat exchange occurs.

It implies that the work done by the gas is positive, and no heat exchange occurs.

Ans: Part A.1

The internal energy of the system has decreased.